inputenc.sty

Alan Jeffrey  Frank Mittelbach

v0.9o 1996/02/14

1 Introduction

This package allows the user to specify an input encoding (for example, ASCII, ISO Latin-1 or Macintosh) by saying:

     \usepackage[encoding name]{inputenc}

The encoding can also be selected in the document with:

     \inputencoding{encoding name}

The encodings provided by this package are:

• ascii ASCII encoding for the range 32–127.

• latin1 ISO Latin-1 encoding.

• latin2 ISO Latin-2 encoding.

• cp850 IBM 850 code page.

• cp437 IBM 437 code page.

• cp437de IBM 437 code page (German version).

• applemac Macintosh encoding.

• next Next encoding.

• ansinew Windows 3.1 ANSI encoding. Each encoding has an associated .def file, for example latin1.def which defines the behaviour of each input character, using the commands:

     \DeclareInputText{slot}{text}
     \DeclareInputMath{slot}{math}

This defines the input character slot to be the text material or math material respectively. For example, latin1.def defines slots "D6 (Æ) and "B5 (μ) by saying:

    \DeclareInputText{"0D6}{\AE}
    \DeclareInputMath{"0B5}{\mu}

Note that the commands should be robust, and should not be dependent on the output encoding. The same slot should not have both a text and a math declaration for it. (This restriction may be removed in future releases of inputenc).

The .def file may also define commands using the declarations:

\providecommand or \ProvideTextCommandDefault,
    \ProvideTextCommandDefault{\textonequarter}{$\frac14$}
    \DeclareInputText{"0BC}{\textonequarter}

This means that a package can change the behaviour of the ‘¡’ slot by redefining the command \textexclamdown. See the documentation in fntguide.tex and ltoutenc.dtx for details of how to declare text commands.

2 Announcing the files

We announce the files:

 1 ⟨package⟩\message{beta test version}
 2 ⟨package⟩\ProvidesPackage{inputenc}
 3 ⟨ascii⟩ \ProvidesFile{ascii.def}
 4 ⟨latin1⟩ \ProvidesFile{latin1.def}
 5 ⟨latin2⟩ \ProvidesFile{latin2.def}
 6 ⟨cp850⟩ \ProvidesFile{cp850.def}
 7 ⟨cp437⟩ \ProvidesFile{cp437.def}
 8 ⟨cp437de⟩ \ProvidesFile{cp437de.def}
 9 ⟨applemac⟩ \ProvidesFile{applemac.def}
10 ⟨next⟩ \ProvidesFile{next.def}
11 ⟨ansinew⟩ \ProvidesFile{ansinew.def}
12 [1996/02/14 v0.9o Input encoding file
13  (test version: liable to change)]
14 \NeedsTeXFormat{LaTeX2e}[1995/12/01]

3 The package

Before we start with the code, an important comment is in order: as you may or may not know, the tabbing environment changes the definition of the commands \', \`, and \=. Outside such an environment these commands produce the corresponding accents, inside they are used for special text positioning, and the accents can be accessed using \a', \a`, and \a=. Therefore we must use the latter instead of the former in the second argument to \DeclareInputText, e.g.

(from latin1.def):
     \DeclareInputText{"0E0}{\@tabacckludge`a}

The command \@tabacckludge is defined (in ltoutenc.dtx) in a way so that \@tabacckludge' will expand to \', so that it is always \' that is carried around internally (the same applies to the other two accent commands).

\DeclareInputText \DeclareInputMath \IeC The declaration commands declare a new active character. The math declaration is the usual trick with \uppercase. The text declaration is sneakier, since in text space matters. We look to see if the definition ends in a macro, by checking whether it’s \meaning ends in a space. If it does, then we add an irrelevant \IeC and braces around the definition, in order to avoid any space after the active char being gobbled up once the text is written out to an auxiliary file.

15 ⟨∗package⟩

16 \def\DeclareInputMath#1{%
17 \bgroup
18  \uccode`\~#1%
19  \uppercase{%
20 \egroup
21  \def~%
22 }%

23 }

24 \def\DeclareInputText#1#2{%
25 \def\@tempa{#2}%
26 \def\@tempb##1 ${}%
27 \edef\@tempa{\expandafter\@tempb\meaning\@tempa$ $}%
28 \ifx\@tempa\@empty
29  \DeclareInputMath{#1}{#2}%
30 \else
31  \DeclareInputMath{#1}{\IeC{#2}}%
32 \fi

33 }

The definition if \IeC is modified not to insert a \protect unless it is needed, this means it works in \hyphenation commands, and other such delicate places.

34 %\DeclareRobustCommand{\IeC}[1]{#1}
35 \def\IeC{%
36 \ifx\protect\@typeset@protect
37  \expandafter\@firstofone
38 \else
39  \protect\IeC
40 \fi}

\inputencoding To set the encoding to be foo, we input an encoding file foo.def with the command \inputencoding. This sets all the slots 128–255 to be active characters, sets their initial definition to be \@undefined, then inputs foo.def.

41 \def\inputencoding#1{%
42 \@tempcnta=128
43 \loop\ifnum\@tempcnta<256
44  \catcode\@tempcnta\active
45  \bgroup
46  \uccode`\~\@tempcnta
47  \uppercase{%
48  \egroup
49  \let~\@undefined
50  }%

51  \advance\@tempcnta\@ne
52 \repeat
53 \input{#1.def}%

54 }

Then for each option, we input that encoding file.

55 \DeclareOption*{\inputencoding{\CurrentOption}}
56 \ProcessOptions

57 ⟨/package⟩

4 Default definitions for characters

Some input characters map to internal function which aren’t found in either T1 or OT1 encoding. For this reason we are providing default definitions for them which will apply unless some output encoding is used which supports those glyphs.

58 ⟨∗latin1 | cp850 | ansinew | next⟩

59 \ProvideTextCommandDefault{\textonequarter}{\ensuremath{\frac14}}
60 \ProvideTextCommandDefault{\textonehalf}{\ensuremath{\frac12}}
61 \ProvideTextCommandDefault{\textthreequarters}{\ensuremath{\frac34}}

62 ⟨/latin1 | cp850 | ansinew | next⟩ 63 ⟨∗applemac | cp850 | ansinew | next⟩

64 \ProvideTextCommandDefault{\textflorin}{\textit{f}}

65 ⟨/applemac | cp850 | ansinew | next⟩ 66 ⟨∗next⟩

67 \ProvideTextCommandDefault{\textquotesinglbase}{,}
68 \ProvideTextCommandDefault{\textquotedblbase}{,,}

69 ⟨/next⟩

Some commands can’t be faked, so we have them generate an error message.

70 ⟨∗latin1 | cp850 | cp437 | cp437de | applemac | ansinew⟩

71 \ProvideTextCommandDefault{\textcent}
72 {\TextSymbolUnavailable\textcent}
73 \ProvideTextCommandDefault{\textyen}
74 {\TextSymbolUnavailable\textyen}

75 ⟨/latin1 | cp850 | cp437 | cp437de | applemac | ansinew⟩ 76 ⟨∗latin1 | latin2 | cp850 | applemac | ansinew⟩

77 \ProvideTextCommandDefault{\textcurrency}
78 {\TextSymbolUnavailable\textcurrency}

79 ⟨/latin1 | latin2 | cp850 | applemac | ansinew⟩ 80 ⟨∗latin1 | cp850 | cp437 | cp437de | ansinew⟩

81 \ProvideTextCommandDefault{\textbrokenbar}
82 {\TextSymbolUnavailable\textbrokenbar}

83 ⟨/latin1 | cp850 | cp437 | cp437de | ansinew⟩ 84 ⟨∗next⟩

85 \ProvideTextCommandDefault{\textquotesingle}
86 {\TextSymbolUnavailable\textquotesingle}

87 ⟨/next⟩

Characters that are supposed to be used only in math will be defined by \providecommand because the L^{A}TEX 2ε assumes that the font encoding for math fonts is static.

88 ⟨∗latin1 | cp850 | cp437 | cp437de | applemac | ansinew | next⟩

 89 \providecommand{\mathdegree}{{^\circ}}

90 ⟨/latin1 | cp850 | cp437 | cp437de | applemac | ansinew | next⟩ 91 ⟨∗latin1 | cp850 | ansinew | next⟩

 92 \providecommand{\mathonesuperior}{{^1}}

93 ⟨/latin1 | cp850 | ansinew | next⟩ 94 ⟨∗latin1 | cp850 | cp437 | cp437de | ansinew | next⟩

 95 \providecommand{\mathtwosuperior}{{^2}}
 96 \providecommand{\maththreesuperior}{{^3}}

97 ⟨/latin1 | cp850 | cp437 | cp437de | ansinew | next⟩ 98 ⟨∗latin1 | applemac | cp850 | cp437 | cp437de | ansinew | next⟩

 99 \providecommand{\mathordmasculine}{{^o}}
100 \providecommand{\mathordfeminine}{{^a}}

101 ⟨/latin1 | applemac | cp850 | cp437 | cp437de | ansinew | next⟩

5 The ASCII encoding

The ASCII encoding only allows characters in the range 32–127, so we only need to provide an empty .def file.

6 The ISO Latin-1 encoding

The ISO Latin-1 encoding file defines the characters in the ISO encoding.

102 ⟨∗latin1⟩

103 \DeclareInputText{"0A0}{\nobreakspace}
104 \DeclareInputMath{"0B0}{\mathdegree}
105 \DeclareInputText{"0A1}{\textexclamdown}
106 \DeclareInputMath{"0B1}{\pm}
107 \DeclareInputText{"0A2}{\textcent}
108 \DeclareInputMath{"0B2}{\mathtwosuperior}
109 \DeclareInputText{"0A3}{\pounds}
110 \DeclareInputMath{"0B3}{\maththreesuperior}
111 \DeclareInputText{"0A4}{\textcurrency}
112 \DeclareInputText{"0B4}{\@tabacckludge'{}}
113 \DeclareInputText{"0A5}{\textyen}
114 \DeclareInputMath{"0B5}{\mu}
115 \DeclareInputText{"0A6}{\textbrokenbar}
116 \DeclareInputText{"0B6}{\P}
117 \DeclareInputText{"0A7}{\S}
118 \DeclareInputText{"0B7}{\textperiodcentered}
119 \DeclareInputText{"0A8}{\"{}}
120 \DeclareInputText{"0B8}{\c\ }
121 \DeclareInputText{"0A9}{\copyright}
122 \DeclareInputMath{"0B9}{\mathonesuperior}

123 \DeclareInputMath{"0AA}{\mathordfeminine}
124 \DeclareInputMath{"0BA}{\mathordmasculine}
125 \DeclareInputText{"0AB}{\guillemotleft}
126 \DeclareInputText{"0BB}{\guillemotright}
127 \DeclareInputMath{"0AC}{\lnot}
128 \DeclareInputText{"0BC}{\textonequarter}
129 \DeclareInputText{"0AD}{\-}
130 \DeclareInputText{"0BD}{\textonehalf}
131 \DeclareInputText{"0AE}{\textregistered}
132 \DeclareInputText{"0BE}{\textthreequarters}
133 \DeclareInputText{"0AF}{\@tabacckludge={}}
134 \DeclareInputText{"0BF}{\textquestiondown}
135 \DeclareInputText{"0C0}{\@tabacckludge`A}
136 \DeclareInputText{"0D0}{\DH}
137 \DeclareInputText{"0C1}{\@tabacckludge'A}
138 \DeclareInputText{"0D1}{\~N}
139 \DeclareInputText{"0C2}{\^A}
140 \DeclareInputText{"0D2}{\@tabacckludge`O}
141 \DeclareInputText{"0C3}{\~A}
142 \DeclareInputText{"0D3}{\@tabacckludge'O}
143 \DeclareInputText{"0C4}{\"A}
144 \DeclareInputText{"0D4}{\^O}
145 \DeclareInputText{"0C5}{\AA}
146 \DeclareInputText{"0D5}{\~O}
147 \DeclareInputText{"0C6}{\AE}
148 \DeclareInputText{"0D6}{\"O}
149 \DeclareInputText{"0C7}{\c C}
150 \DeclareInputMath{"0D7}{\times}
151 \DeclareInputText{"0C8}{\@tabacckludge`E}
152 \DeclareInputText{"0D8}{\O}
153 \DeclareInputText{"0C9}{\@tabacckludge'E}
154 \DeclareInputText{"0D9}{\@tabacckludge`U}
155 \DeclareInputText{"0CA}{\^E}
156 \DeclareInputText{"0DA}{\@tabacckludge'U}
157 \DeclareInputText{"0CB}{\"E}
158 \DeclareInputText{"0DB}{\^U}
159 \DeclareInputText{"0CC}{\@tabacckludge`I}
160 \DeclareInputText{"0DC}{\"U}
161 \DeclareInputText{"0CD}{\@tabacckludge'I}
162 \DeclareInputText{"0DD}{\@tabacckludge'Y}
163 \DeclareInputText{"0CE}{\^I}
164 \DeclareInputText{"0DE}{\TH}
165 \DeclareInputText{"0CF}{\"I}
166 \DeclareInputText{"0DF}{\ss}
167 \DeclareInputText{"0E0}{\@tabacckludge`a}
168 \DeclareInputText{"0F0}{\dh}
169 \DeclareInputText{"0E1}{\@tabacckludge'a}
170 \DeclareInputText{"0F1}{\~n}
171 \DeclareInputText{"0E2}{\^a}

172 \DeclareInputText{"0F2}{\@tabacckludge`o}
173 \DeclareInputText{"0E3}{\~a}
174 \DeclareInputText{"0F3}{\@tabacckludge'o}
175 \DeclareInputText{"0E4}{\"a}
176 \DeclareInputText{"0F4}{\^o}
177 \DeclareInputText{"0E5}{\aa}
178 \DeclareInputText{"0F5}{\~o}
179 \DeclareInputText{"0E6}{\ae}
180 \DeclareInputText{"0F6}{\"o}
181 \DeclareInputText{"0E7}{\c c}
182 \DeclareInputMath{"0F7}{\div}
183 \DeclareInputText{"0E8}{\@tabacckludge`e}
184 \DeclareInputText{"0F8}{\o}
185 \DeclareInputText{"0E9}{\@tabacckludge'e}
186 \DeclareInputText{"0F9}{\@tabacckludge`u}
187 \DeclareInputText{"0EA}{\^e}
188 \DeclareInputText{"0FA}{\@tabacckludge'u}
189 \DeclareInputText{"0EB}{\"e}
190 \DeclareInputText{"0FB}{\^u}
191 \DeclareInputText{"0EC}{\@tabacckludge`\i}
192 \DeclareInputText{"0FC}{\"u}
193 \DeclareInputText{"0ED}{\@tabacckludge'\i}
194 \DeclareInputText{"0FD}{\@tabacckludge'y}
195 \DeclareInputText{"0EE}{\^\i}
196 \DeclareInputText{"0FE}{\th}
197 \DeclareInputText{"0EF}{\"\i}
198 \DeclareInputText{"0FF}{\"y}

199 ⟨/latin1⟩

7 The ISO Latin-2 encoding

The ISO Latin-2 encoding file defines the characters in the ISO 8859-2 encoding. It was contributed by Petr Sojka (sojka@muni.cz) with slight technical updates by Frank.

Then the characters themselves:

200 ⟨∗latin2⟩

201 \DeclareInputText{"0A0}{\nobreakspace}
202 \DeclareInputText{"0B0}{\r{}}
203 \DeclareInputText{"0A1}{\k A}
204 \DeclareInputText{"0B1}{\k a}
205 \DeclareInputText{"0A2}{\u{}}
206 \DeclareInputText{"0B2}{\k\ }
207 \DeclareInputText{"0A3}{\L}
208 \DeclareInputText{"0B3}{\l}
209 \DeclareInputText{"0A4}{\textcurrency}
210 \DeclareInputText{"0B4}{\@tabacckludge'{}}
211 \DeclareInputText{"0A5}{\v L}
212 \DeclareInputText{"0B5}{\v l}

213 \DeclareInputText{"0A6}{\@tabacckludge'S}
214 \DeclareInputText{"0B6}{\@tabacckludge's}
215 \DeclareInputText{"0A7}{\S}
216 \DeclareInputText{"0B7}{\v{}}
217 \DeclareInputText{"0A8}{\"{}}
218 \DeclareInputText{"0B8}{\c\ }
219 \DeclareInputText{"0A9}{\v S}
220 \DeclareInputText{"0B9}{\v s}
221 \DeclareInputText{"0AA}{\c S}
222 \DeclareInputText{"0BA}{\c s}
223 \DeclareInputText{"0AB}{\v T}
224 \DeclareInputText{"0BB}{\v t}
225 \DeclareInputText{"0AC}{\@tabacckludge'Z}
226 \DeclareInputText{"0BC}{\@tabacckludge'z}
227 \DeclareInputText{"0AD}{\-}
228 \DeclareInputText{"0BD}{\H{}}
229 \DeclareInputText{"0AE}{\v Z}
230 \DeclareInputText{"0BE}{\v z}
231 \DeclareInputText{"0AF}{\.Z}
232 \DeclareInputText{"0BF}{\.z}
233 \DeclareInputText{"0C0}{\@tabacckludge'R}
234 \DeclareInputText{"0D0}{\DH}
235 \DeclareInputText{"0C1}{\@tabacckludge'A}
236 \DeclareInputText{"0D1}{\@tabacckludge'N}
237 \DeclareInputText{"0C2}{\^A}
238 \DeclareInputText{"0D2}{\v N}
239 \DeclareInputText{"0C3}{\u A}
240 \DeclareInputText{"0D3}{\@tabacckludge'O}
241 \DeclareInputText{"0C4}{\"A}
242 \DeclareInputText{"0D4}{\^O}
243 \DeclareInputText{"0C5}{\@tabacckludge'L}
244 \DeclareInputText{"0D5}{\H O}
245 \DeclareInputText{"0C6}{\@tabacckludge'C}
246 \DeclareInputText{"0D6}{\"O}
247 \DeclareInputText{"0C7}{\c C}
248 \DeclareInputMath{"0D7}{\times}
249 \DeclareInputText{"0C8}{\v C}
250 \DeclareInputText{"0D8}{\v R}
251 \DeclareInputText{"0C9}{\@tabacckludge'E}
252 \DeclareInputText{"0D9}{\r U}
253 \DeclareInputText{"0CA}{\k E}
254 \DeclareInputText{"0DA}{\@tabacckludge'U}
255 \DeclareInputText{"0CB}{\"E}
256 \DeclareInputText{"0DB}{\H U}
257 \DeclareInputText{"0CC}{\v E}
258 \DeclareInputText{"0DC}{\"U}
259 \DeclareInputText{"0CD}{\@tabacckludge'I}
260 \DeclareInputText{"0DD}{\@tabacckludge'Y}
261 \DeclareInputText{"0CE}{\^I}

262 \DeclareInputText{"0DE}{\c T}
263 \DeclareInputText{"0CF}{\v D}
264 \DeclareInputText{"0DF}{\ss}
265 \DeclareInputText{"0E0}{\@tabacckludge'r}
266 \DeclareInputText{"0F0}{\dh}
267 \DeclareInputText{"0E1}{\@tabacckludge'a}
268 \DeclareInputText{"0F1}{\@tabacckludge'n}
269 \DeclareInputText{"0E2}{\^a}
270 \DeclareInputText{"0F2}{\v n}
271 \DeclareInputText{"0E3}{\u a}
272 \DeclareInputText{"0F3}{\@tabacckludge'o}
273 \DeclareInputText{"0E4}{\"a}
274 \DeclareInputText{"0F4}{\^o}
275 \DeclareInputText{"0E5}{\@tabacckludge'l}
276 \DeclareInputText{"0F5}{\H o}
277 \DeclareInputText{"0E6}{\@tabacckludge'c}
278 \DeclareInputText{"0F6}{\"o}
279 \DeclareInputText{"0E7}{\c c}
280 \DeclareInputMath{"0F7}{\div}
281 \DeclareInputText{"0E8}{\v c}
282 \DeclareInputText{"0F8}{\v r}
283 \DeclareInputText{"0E9}{\@tabacckludge'e}
284 \DeclareInputText{"0F9}{\r u}
285 \DeclareInputText{"0EA}{\k e}
286 \DeclareInputText{"0FA}{\@tabacckludge'u}
287 \DeclareInputText{"0EB}{\"e}
288 \DeclareInputText{"0FB}{\H u}
289 \DeclareInputText{"0EC}{\v e}
290 \DeclareInputText{"0FC}{\"u}
291 \DeclareInputText{"0ED}{\@tabacckludge'\i}
292 \DeclareInputText{"0FD}{\@tabacckludge'y}
293 \DeclareInputText{"0EE}{\^\i}
294 \DeclareInputText{"0FE}{\c t}
295 \DeclareInputText{"0EF}{\v d}
296 \DeclareInputText{"0FF}{\.{}}

297 ⟨/latin2⟩

8 The IBM code page 850

The coding for this input encoding was contributed by Timo Knuutila (knuutila@

 cs.utu.fi), and edited by Christian Bartels (ii140ba@vm1.rz.rwth-aachen.de).

The dos graphics ‘letters’ and a few other positions are ignored (left undefined).

298 ⟨∗cp850⟩

299 \DeclareInputText{"080}{\c C}
300 \DeclareInputText{"090}{\@tabacckludge'E}
301 \DeclareInputText{"081}{\"u}
302 \DeclareInputText{"091}{\ae}
303 \DeclareInputText{"082}{\@tabacckludge'e}

304 \DeclareInputText{"092}{\AE}
305 \DeclareInputText{"083}{\^a}
306 \DeclareInputText{"093}{\^o}
307 \DeclareInputText{"084}{\"a}
308 \DeclareInputText{"094}{\"o}
309 \DeclareInputText{"085}{\@tabacckludge`a}
310 \DeclareInputText{"095}{\@tabacckludge`o}
311 \DeclareInputText{"086}{\aa}
312 \DeclareInputText{"096}{\^u}
313 \DeclareInputText{"087}{\c c}
314 \DeclareInputText{"097}{\@tabacckludge`u}
315 \DeclareInputText{"088}{\^e}
316 \DeclareInputText{"098}{\"y}
317 \DeclareInputText{"089}{\"e}
318 \DeclareInputText{"099}{\"O}
319 \DeclareInputText{"08A}{\@tabacckludge`e}
320 \DeclareInputText{"09A}{\"U}
321 \DeclareInputText{"08B}{\"\i}
322 \DeclareInputText{"09B}{\o}
323 \DeclareInputText{"08C}{\^\i}
324 \DeclareInputText{"09C}{\pounds}
325 \DeclareInputText{"08D}{\@tabacckludge`\i}
326 \DeclareInputText{"09D}{\O}
327 \DeclareInputText{"08E}{\"A}
328 \DeclareInputMath{"09E}{\times}
329 \DeclareInputText{"08F}{\AA}
330 \DeclareInputText{"09F}{\textflorin}
331 \DeclareInputText{"0A0}{\@tabacckludge'a}
332 \DeclareInputText{"0A1}{\@tabacckludge'\i}
333 \DeclareInputText{"0A2}{\@tabacckludge'o}
334 \DeclareInputText{"0A3}{\@tabacckludge'u}
335 \DeclareInputText{"0A4}{\~n}               %% "B0-"B4: DG
336 \DeclareInputText{"0A5}{\~N}
337 \DeclareInputText{"0B5}{\@tabacckludge'A}
338 \DeclareInputMath{"0A6}{\mathordmasculine}
339 \DeclareInputText{"0B6}{\^A}
340 \DeclareInputMath{"0A7}{\mathordfeminine}
341 \DeclareInputText{"0B7}{\@tabacckludge`A}
342 \DeclareInputText{"0A8}{\textquestiondown}
343 \DeclareInputText{"0B8}{\copyright}
344 \DeclareInputText{"0A9}{\textregistered}
345 \DeclareInputMath{"0AA}{\lnot}
346 \DeclareInputText{"0AB}{\textonehalf}
347 \DeclareInputText{"0AC}{\textonequarter}   %% "B9-"BC: DG
348 \DeclareInputText{"0AD}{\textexclamdown}
349 \DeclareInputText{"0BD}{\textcent}
350 \DeclareInputText{"0AE}{\guillemotleft}
351 \DeclareInputText{"0BE}{\textyen}
352 \DeclareInputText{"0AF}{\guillemotright}

353 % "BF-"C5: DG
354 \DeclareInputText{"0D0}{\dh}
355 \DeclareInputText{"0D1}{\DH}
356 \DeclareInputText{"0D2}{\^E}
357 \DeclareInputText{"0D3}{\"E}
358 \DeclareInputText{"0D4}{\@tabacckludge`E}
359 \DeclareInputText{"0D5}{\i}
360 \DeclareInputText{"0C6}{\~a}
361 \DeclareInputText{"0D6}{\@tabacckludge'I}
362 \DeclareInputText{"0C7}{\~A}
363 \DeclareInputText{"0D7}{\^I}
364 \DeclareInputText{"0D8}{\"I}
365 % "C8-"CE: DG
366 % "D9-"DC: DG
367 \DeclareInputText{"0DD}{\textbrokenbar}
368 \DeclareInputText{"0DE}{\@tabacckludge`I}
369 \DeclareInputText{"0CF}{\textcurrency}  %% "DF: DG
370 \DeclareInputText{"0E0}{\@tabacckludge'O}
371 \DeclareInputText{"0F0}{\textendash}
372 \DeclareInputText{"0E1}{\ss}
373 \DeclareInputMath{"0F1}{\pm}
374 \DeclareInputText{"0E2}{\^O}   %% "F2: DG (double underline)
375 \DeclareInputText{"0E3}{\@tabacckludge`O}
376 \DeclareInputText{"0F3}{\textthreequarters}
377 \DeclareInputText{"0E4}{\~o}
378 \DeclareInputText{"0F4}{\P}
379 \DeclareInputText{"0E5}{\~O}
380 \DeclareInputText{"0F5}{\S}
381 \DeclareInputMath{"0E6}{\mu}
382 \DeclareInputMath{"0F6}{\div}
383 \DeclareInputText{"0E7}{\th}
384 \DeclareInputText{"0F7}{\c\ }
385 \DeclareInputText{"0E8}{\TH}
386 \DeclareInputMath{"0F8}{\mathdegree}
387 \DeclareInputText{"0E9}{\@tabacckludge'U}
388 \DeclareInputText{"0F9}{\"{}}
389 \DeclareInputText{"0EA}{\^U}
390 \DeclareInputText{"0FA}{\textperiodcentered}
391 \DeclareInputText{"0EB}{\@tabacckludge`U}
392 \DeclareInputMath{"0FB}{\mathonesuperior}
393 \DeclareInputText{"0EC}{\@tabacckludge'y}
394 \DeclareInputMath{"0FC}{\maththreesuperior}
395 \DeclareInputText{"0ED}{\@tabacckludge'Y}
396 \DeclareInputMath{"0FD}{\mathtwosuperior}
397 \DeclareInputText{"0EE}{\@tabacckludge={}}  %% "FE: DG (box)
398 \DeclareInputText{"0EF}{\@tabacckludge'{}}
399 \DeclareInputText{"0FF}{\nobreakspace}

400 ⟨/cp850⟩

9 The IBM codepage 437

The coding for this inputencoding is based on work by Volker Kunert (volker@

 numsun1.mathematik.uni-halle.de) and bontus@al6000.physik.uni-siegen.de.

The dos graphics ‘letters’ and a few other positions are ignored (left undefined).

Unfortunately there is no agreement as to whether slot E1 should be ‘β’ or ‘ß’, so we provide two variants, one (cp437) with ‘β’ and one (cp437de) with ‘ß’.

401 ⟨∗cp437 | cp437de⟩

402 \DeclareInputText{"080}{{\c C}}
403 \DeclareInputText{"081}{\"u}
404 \DeclareInputText{"082}{\@tabacckludge'e}
405 \DeclareInputText{"083}{\^a}
406 \DeclareInputText{"084}{\"a}
407 \DeclareInputText{"085}{\@tabacckludge`a}
408 \DeclareInputText{"086}{\aa}
409 \DeclareInputText{"087}{{\c c}}
410 \DeclareInputText{"088}{\^e}
411 \DeclareInputText{"089}{\"e}
412 \DeclareInputText{"08A}{\@tabacckludge`e}
413 \DeclareInputText{"08B}{\"\i}
414 \DeclareInputText{"08C}{\^\i}
415 \DeclareInputText{"08D}{\@tabacckludge`\i}
416 \DeclareInputText{"08E}{\"A}
417 \DeclareInputText{"08F}{\AA}
418 \DeclareInputText{"090}{\@tabacckludge'E}
419 \DeclareInputText{"091}{\ae}
420 \DeclareInputText{"092}{\AE}
421 \DeclareInputText{"093}{\^o}
422 \DeclareInputText{"094}{\"o}
423 \DeclareInputText{"095}{\@tabacckludge`o}
424 \DeclareInputText{"096}{\^u}
425 \DeclareInputText{"097}{\@tabacckludge`u}
426 \DeclareInputText{"098}{\"y}
427 \DeclareInputText{"099}{\"O}
428 \DeclareInputText{"09A}{\"U}
429 \DeclareInputText{"09B}{\textcent}
430 \DeclareInputText{"09C}{\pounds}
431 \DeclareInputText{"09D}{\textyen}
432 %\DeclareInputMath{"09E}{} % Pt
433 % should be int with dash
434 %\DeclareInputMath{"09F}{{\ensuremath{\rlap{-}\int}}}
435 \DeclareInputText{"0A0}{\@tabacckludge'a} % 160
436 \DeclareInputText{"0A1}{\@tabacckludge'\i}
437 \DeclareInputText{"0A2}{\@tabacckludge'o}
438 \DeclareInputText{"0A3}{\@tabacckludge'u}
439 \DeclareInputText{"0A4}{\~n}
440 \DeclareInputText{"0A5}{\~N}
441 \DeclareInputMath{"0A6}{\mathordfeminine}
442 \DeclareInputMath{"0A7}{\mathordmasculine}

443 \DeclareInputText{"0A8}{\textquestiondown}
444 %\DeclareInputText{"0A9}{\ensuremath{\lceil}} % left upper corner
445 %\DeclareInputMath{"0AA}{\ensuremath{\rceil}} % right upper corner
446 \DeclareInputText{"0AB}{\textonehalf}
447 \DeclareInputText{"0AC}{\textonequarter}
448 \DeclareInputText{"0AD}{\textexclamdown}
449 \DeclareInputText{"0AE}{\guillemotleft}
450 \DeclareInputText{"0AF}{\guillemotright}
451 %\DeclareInputText{"0B0}{\textlightgraybox}
452 %\DeclareInputText{"0B1}{\textgraybox}
453 %\DeclareInputText{"0B2}{\textdarkgraybox}
454 %\DeclareInputMath{"0B3}{} % vertical bar
455 %\DeclareInputText{"0B4}{} % vertical bar with branch to left
456 %\DeclareInputText{"0B5}{} % vertical bar with double branch to left
457 %\DeclareInputText{"0B6}{} % double bar with single branch to left
458 %\DeclareInputText{"0B7}{} % graphic
459 %\DeclareInputText{"0B8}{} % graphic
460 %\DeclareInputMath{"0B9}{} % vertical double bar with branch to left
461 %\DeclareInputMath{"0BA}{} % vertical double bar
462 %\DeclareInputText{"0BB}{} % double upper right corner
463 %\DeclareInputMath{"0BC}{} % double lower right corner
464 %\DeclareInputText{"0BD}{} % graphic
465 %\DeclareInputText{"0BE}{} % graphic
466 %\DeclareInputMath{"0BF}{\ensuremath{\rceil}}
467 %\DeclareInputMath{"0C0}{\ensuremath{\lfloor}}
468 %\DeclareInputText{"0C1}{} % dash with branch up
469 %\DeclareInputText{"0C2}{} % dash with branch down
470 %\DeclareInputText{"0C3}{} % vertical bar with branch to right
471 \DeclareInputText{"0C4}{\textendash}
472 %\DeclareInputText{"0C5}{} % vertical bar crossed with dash
473 %\DeclareInputText{"0C6}{} % graphic
474 %\DeclareInputText{"0C7}{} % graphic
475 %\DeclareInputText{"0C8}{} % double lower left corner
476 %\DeclareInputText{"0C9}{} % double upper left corner
477 %\DeclareInputText{"0CA}{} % double dash with branch up
478 %\DeclareInputText{"0CB}{} % double dash with branch down
479 %\DeclareInputText{"0CC}{} % double bar with branch right
480 %\DeclareInputText{"0CD}{=} % double dash
481 %\DeclareInputText{"0CE}{} % double bar crossing double dash
482 %\DeclareInputText{"0CF}{} % graphic
483 %\DeclareInputMath{"0D0}{}
484 %\DeclareInputText{"0D1}{}
485 %\DeclareInputText{"0D2}{}
486 %\DeclareInputText{"0D3}{}
487 %\DeclareInputText{"0D4}{}
488 %\DeclareInputText{"0D5}{}
489 %\DeclareInputText{"0D6}{}
490 %\DeclareInputText{"0D7}{}
491 %\DeclareInputText{"0D8}{}
492 %\DeclareInputMath{"0D9}{\ensuremath{\rfloor}} % lower right corner

493 %\DeclareInputMath{"0DA}{\ensuremath{\lceil}}  % upper left corner
494 %\DeclareInputText{"0DB}{} % black box
495 %\DeclareInputText{"0DC}{} % lower half of black box
496 \DeclareInputText{"0DD}{\textbrokenbar} % bar with gap
497 %\DeclareInputText{"0DE}{}
498 %\DeclareInputText{"0DF}{} % upper half of black box
499 \DeclareInputMath{"0E0}{\alpha}

500 ⟨/cp437 | cp437de⟩

501 ⟨cp437⟩\DeclareInputMath{"0E1}{\beta}
502 ⟨cp437de⟩\DeclareInputText{"0E1}{\ss}

503 ⟨∗cp437 | cp437de⟩

504 \DeclareInputMath{"0E2}{\Gamma}
505 \DeclareInputMath{"0E3}{\Pi}
506 \DeclareInputMath{"0E4}{\Sigma}
507 \DeclareInputMath{"0E5}{\sigma}
508 \DeclareInputMath{"0E6}{\mu}
509 \DeclareInputMath{"0E7}{\gamma}
510 \DeclareInputMath{"0E8}{\Theta}
511 \DeclareInputMath{"0E9}{\theta}
512 \DeclareInputMath{"0EA}{\Omega}
513 \DeclareInputMath{"0EB}{\delta}
514 \DeclareInputMath{"0EC}{\infty}
515 \DeclareInputMath{"0ED}{\emptyset}
516 \DeclareInputMath{"0EE}{\varepsilon}
517 \DeclareInputMath{"0EF}{\cap}
518 \DeclareInputMath{"0F0}{\equiv}
519 \DeclareInputMath{"0F1}{\pm}
520 \DeclareInputMath{"0F2}{\ge}
521 \DeclareInputMath{"0F3}{\le}
522 %\DeclareInputMath{"0F4}{}   % upper part of integral sign
523 %\DeclareInputMath{"0F5}{}   % lower part of integral sign
524 \DeclareInputMath{"0F6}{\div}
525 \DeclareInputText{"0F7}{\c\ } % cedilla ?
526 \DeclareInputMath{"0F8}{\mathdegree}
527 \DeclareInputText{"0F9}{\textperiodcentered}
528 \DeclareInputText{"0FA}{\textbullet}
529 \DeclareInputMath{"0FB}{\surd}
530 \DeclareInputMath{"0FC}{\maththreesuperior}
531 \DeclareInputMath{"0FD}{\mathtwosuperior}
532 %\DeclareInputText{"0FE}{} % black square
533 \DeclareInputText{"0FF}{\nobreakspace}

534 ⟨/cp437 | cp437de⟩

10 The Macintosh encoding

The coding for this input encoding was contributed by Constantin Kahn (kahn@ math.toronto.edu), with minor modifications by Alan Jeffrey.

535 ⟨∗applemac⟩

536 \DeclareInputText{"080}{\"A}
537 \DeclareInputText{"081}{\AA}
538 \DeclareInputText{"082}{\c C}
539 \DeclareInputText{"083}{\@tabacckludge'E}
540 \DeclareInputText{"084}{\~N}
541 \DeclareInputText{"085}{\"O}
542 \DeclareInputText{"086}{\"U}
543 \DeclareInputText{"087}{\@tabacckludge'a}
544 \DeclareInputText{"088}{\@tabacckludge`a}
545 \DeclareInputText{"089}{\^a}
546 \DeclareInputText{"08A}{\"a}
547 \DeclareInputText{"08B}{\~a}
548 \DeclareInputText{"08C}{\aa}
549 \DeclareInputText{"08D}{\c c}
550 \DeclareInputText{"08E}{\@tabacckludge'e}
551 \DeclareInputText{"08F}{\@tabacckludge`e}
552 \DeclareInputText{"090}{\^e}
553 \DeclareInputText{"091}{\"e}
554 \DeclareInputText{"092}{\@tabacckludge'\i}
555 \DeclareInputText{"093}{\@tabacckludge`\i}
556 \DeclareInputText{"094}{\^\i}
557 \DeclareInputText{"095}{\"\i}
558 \DeclareInputText{"096}{\~n}
559 \DeclareInputText{"097}{\@tabacckludge'o}
560 \DeclareInputText{"098}{\@tabacckludge`o}
561 \DeclareInputText{"099}{\^o}
562 \DeclareInputText{"09A}{\"o}
563 \DeclareInputText{"09B}{\~o}
564 \DeclareInputText{"09C}{\@tabacckludge'u}
565 \DeclareInputText{"09D}{\@tabacckludge`u}
566 \DeclareInputText{"09E}{\^u}
567 \DeclareInputText{"09F}{\"u}
568 \DeclareInputText{"0A0}{\dag}
569 \DeclareInputMath{"0A1}{\mathdegree}
570 \DeclareInputText{"0A2}{\textcent}
571 \DeclareInputText{"0A3}{\pounds}
572 \DeclareInputText{"0A4}{\S}
573 \DeclareInputText{"0A5}{\textbullet}
574 \DeclareInputText{"0A6}{\P}
575 \DeclareInputText{"0A7}{\ss}
576 \DeclareInputText{"0A8}{\textregistered}
577 \DeclareInputText{"0A9}{\copyright}
578 \DeclareInputText{"0AA}{\texttrademark}
579 \DeclareInputText{"0AB}{\@tabacckludge'{}}
580 \DeclareInputText{"0AC}{\"{}}
581 \DeclareInputMath{"0AD}{\neq}
582 \DeclareInputText{"0AE}{\AE}
583 \DeclareInputText{"0AF}{\O}
584 \DeclareInputMath{"0B0}{\infty}
585 \DeclareInputMath{"0B1}{\pm}

586 \DeclareInputMath{"0B2}{\leq}
587 \DeclareInputMath{"0B3}{\geq}
588 \DeclareInputText{"0B4}{\textyen}
589 \DeclareInputMath{"0B5}{\mu}
590 \DeclareInputMath{"0B6}{\partial}
591 \DeclareInputMath{"0B7}{\Sigma}
592 \DeclareInputMath{"0B8}{\Pi}
593 \DeclareInputMath{"0B9}{\pi}
594 \DeclareInputMath{"0BA}{\int}
595 \DeclareInputMath{"0BB}{\mathordfeminine}
596 \DeclareInputMath{"0BC}{\mathordmasculine}
597 \DeclareInputMath{"0BD}{\Omega}
598 \DeclareInputText{"0BE}{\ae}
599 \DeclareInputText{"0BF}{\o}
600 \DeclareInputText{"0C0}{\textquestiondown}
601 \DeclareInputText{"0C1}{\textexclamdown}
602 \DeclareInputMath{"0C2}{\lnot}
603 \DeclareInputMath{"0C3}{\surd}
604 \DeclareInputText{"0C4}{\textflorin}
605 \DeclareInputMath{"0C5}{\approx}
606 \DeclareInputMath{"0C6}{\Delta}
607 \DeclareInputText{"0C7}{\guillemotleft}
608 \DeclareInputText{"0C8}{\guillemotright}
609 \DeclareInputText{"0C9}{\dots}
610 \DeclareInputText{"0CA}{\nobreakspace}
611 \DeclareInputText{"0CB}{\@tabacckludge`A}
612 \DeclareInputText{"0CC}{\~A}
613 \DeclareInputText{"0CD}{\~O}
614 \DeclareInputText{"0CE}{\OE}
615 \DeclareInputText{"0CF}{\oe}
616 \DeclareInputText{"0D0}{\textendash}
617 \DeclareInputText{"0D1}{\textemdash}
618 \DeclareInputText{"0D2}{\textquotedblleft}
619 \DeclareInputText{"0D3}{\textquotedblright}
620 \DeclareInputText{"0D4}{\textquoteleft}
621 \DeclareInputText{"0D5}{\textquoteright}
622 \DeclareInputMath{"0D6}{\div}
623 \DeclareInputMath{"0D7}{\diamond}
624 \DeclareInputText{"0D8}{\"y}
625 \DeclareInputText{"0D9}{\"Y}
626 \DeclareInputMath{"0DA}{/}
627 \DeclareInputText{"0DB}{\textcurrency}
628 \DeclareInputText{"0DC}{\guilsinglleft}
629 \DeclareInputText{"0DD}{\guilsinglright}
630 \DeclareInputText{"0DE}{fi}
631 \DeclareInputText{"0DF}{fl}
632 \DeclareInputText{"0E0}{\ddag}
633 \DeclareInputText{"0E1}{\textperiodcentered}
634 \DeclareInputText{"0E2}{\quotesinglbase}
635 \DeclareInputText{"0E3}{\quotedblbase}

636 \DeclareInputText{"0E4}{\perthousand}
637 \DeclareInputText{"0E5}{\^A}
638 \DeclareInputText{"0E6}{\^E}
639 \DeclareInputText{"0E7}{\@tabacckludge'A}
640 \DeclareInputText{"0E8}{\"E}
641 \DeclareInputText{"0E9}{\@tabacckludge`E}
642 \DeclareInputText{"0EA}{\@tabacckludge'I}
643 \DeclareInputText{"0EB}{\^I}
644 \DeclareInputText{"0EC}{\"I}
645 \DeclareInputText{"0ED}{\@tabacckludge`I}
646 \DeclareInputText{"0EE}{\@tabacckludge'O}
647 \DeclareInputText{"0EF}{\^O}
648 \DeclareInputText{"0F0}{\textapplelogo}
649 \DeclareInputText{"0F1}{\@tabacckludge`O}
650 \DeclareInputText{"0F2}{\@tabacckludge'U}
651 \DeclareInputText{"0F3}{\^U}
652 \DeclareInputText{"0F4}{\@tabacckludge`U}
653 \DeclareInputText{"0F5}{\i}
654 \DeclareInputText{"0F6}{\^{}}
655 \DeclareInputText{"0F7}{\~{}}
656 \DeclareInputText{"0F8}{\@tabacckludge={}}
657 \DeclareInputText{"0F9}{\u{}}
658 \DeclareInputText{"0FA}{\.{}}
659 \DeclareInputText{"0FB}{\r{}}
660 \DeclareInputText{"0FC}{\c\ }
661 \DeclareInputText{"0FD}{\H{}}
662 \DeclareInputText{"0FE}{\k{}}
663 \DeclareInputText{"0FF}{\v{}}

664 ⟨/applemac⟩

11 The Next encoding

The coding for this input encoding is based on work by Stefan Ried (stef@

 theo-phys.uni-essen.de and Holger Uhr (huhr@uni-paderborn.de).

665 ⟨∗next⟩

666 \DeclareInputText{"080}{\nobreakspace}
667 \DeclareInputText{"081}{\@tabacckludge`A}
668 \DeclareInputText{"082}{\@tabacckludge'A}
669 \DeclareInputText{"083}{\^A}
670 \DeclareInputText{"084}{\~A}
671 \DeclareInputText{"085}{\"A}
672 \DeclareInputText{"086}{\AA}
673 \DeclareInputText{"087}{\c C}
674 \DeclareInputText{"088}{\@tabacckludge`E}
675 \DeclareInputText{"089}{\@tabacckludge'E}
676 \DeclareInputText{"08A}{\^E}
677 \DeclareInputText{"08B}{\"E}
678 \DeclareInputText{"08C}{\@tabacckludge`I}

679 \DeclareInputText{"08D}{\@tabacckludge'I}
680 \DeclareInputText{"08E}{\^I}
681 \DeclareInputText{"08F}{\"I}
682 %\DeclareInputText{"090}{?}
683 \DeclareInputText{"091}{\~N}
684 \DeclareInputText{"092}{\@tabacckludge`O}
685 \DeclareInputText{"093}{\@tabacckludge'O}
686 \DeclareInputText{"094}{\^O}
687 \DeclareInputText{"095}{\~O}
688 \DeclareInputText{"096}{\"O}
689 \DeclareInputText{"097}{\@tabacckludge`U}
690 \DeclareInputText{"098}{\@tabacckludge'U}
691 \DeclareInputText{"099}{\^U}
692 \DeclareInputText{"09A}{\"U}
693 \DeclareInputText{"09B}{\@tabacckludge'Y}
694 %\DeclareInputText{"09C}{?}
695 \DeclareInputMath{"09D}{\mu}
696 \DeclareInputMath{"09E}{\times}
697 \DeclareInputMath{"09F}{\div}
698 \DeclareInputText{"0A0}{\copyright}
699 \DeclareInputText{"0A1}{\textexclamdown}
700 %\DeclareInputText{"0A2}{?}
701 \DeclareInputText{"0A3}{\pounds}
702 \DeclareInputMath{"0A4}{/}
703 %\DeclareInputText{"0A5}{?}
704 \DeclareInputText{"0A6}{\textflorin}
705 \DeclareInputText{"0A7}{\S}
706 %\DeclareInputText{"0A8}{?}
707 \DeclareInputText{"0A9}{\textquotesingle}
708 \DeclareInputText{"0AA}{\textquotedblleft}
709 %\DeclareInputText{"0AB}{?}
710 %\DeclareInputText{"0AC}{?}
711 %\DeclareInputText{"0AD}{?}
712 \DeclareInputText{"0AE}{fi}
713 \DeclareInputText{"0AF}{fl}
714 %\DeclareInputText{"0B0}{?}
715 \DeclareInputText{"0B1}{\textendash}
716 \DeclareInputText{"0B2}{\dag}
717 \DeclareInputText{"0B3}{\ddag}
718 \DeclareInputText{"0B4}{\textperiodcentered}
719 \DeclareInputText{"0B5}{\textbar}
720 \DeclareInputText{"0B6}{\P}
721 \DeclareInputText{"0B7}{\textbullet}
722 \DeclareInputText{"0B8}{\textquotesinglbase}
723 \DeclareInputText{"0B9}{\textquotedblbase}
724 \DeclareInputText{"0BA}{\textquotedblright}
725 %\DeclareInputText{"0BB}{?}
726 \DeclareInputText{"0BC}{\textellipsis}
727 %\DeclareInputText{"0BD}{?}
728 \DeclareInputMath{"0BE}{\lnot}

729 \DeclareInputText{"0BF}{\textquestiondown}
730 \DeclareInputMath{"0C0}{\mathonesuperior}
731 \DeclareInputText{"0C1}{\@tabacckludge`{}}
732 \DeclareInputText{"0C2}{\@tabacckludge'{}}
733 \DeclareInputText{"0C3}{\^{}}
734 \DeclareInputText{"0C4}{\~{}}
735 \DeclareInputText{"0C5}{\@tabacckludge={}}
736 \DeclareInputText{"0C6}{\u{}}
737 \DeclareInputText{"0C7}{\.{}}
738 \DeclareInputText{"0C8}{\"{}}
739 \DeclareInputMath{"0C9}{\mathtwosuperior}
740 \DeclareInputMath{"0CA}{\mathdegree}
741 \DeclareInputText{"0CB}{\c\ }
742 \DeclareInputMath{"0CC}{\maththreesuperior}
743 \DeclareInputText{"0CD}{\H{}}
744 %\DeclareInputText{"0CE}{?}
745 \DeclareInputText{"0CF}{\v{}}
746 \DeclareInputText{"0D0}{\textemdash}
747 \DeclareInputMath{"0D1}{\pm}
748 \DeclareInputText{"0D2}{\textonequarter}
749 \DeclareInputText{"0D3}{\textonehalf}
750 \DeclareInputText{"0D4}{\textthreequarters}
751 \DeclareInputText{"0D5}{\@tabacckludge`a}
752 \DeclareInputText{"0D6}{\@tabacckludge'a}
753 \DeclareInputText{"0D7}{\^a}
754 \DeclareInputText{"0D8}{\~a}
755 \DeclareInputText{"0D9}{\"a}
756 \DeclareInputText{"0DA}{\aa}
757 \DeclareInputText{"0DB}{\c c}
758 \DeclareInputText{"0DC}{\@tabacckludge`e}
759 \DeclareInputText{"0DD}{\@tabacckludge'e}
760 \DeclareInputText{"0DE}{\^e}
761 \DeclareInputText{"0DF}{\"e}
762 \DeclareInputText{"0E0}{\@tabacckludge`\i}
763 \DeclareInputText{"0E1}{\AE}
764 \DeclareInputText{"0E2}{\@tabacckludge'\i}
765 \DeclareInputMath{"0E3}{\mathordfeminine}
766 \DeclareInputText{"0E4}{\^\i}
767 \DeclareInputText{"0E5}{\"\i}
768 %\DeclareInputText{"0E6}{?}
769 \DeclareInputText{"0E7}{\~n}
770 \DeclareInputText{"0E8}{\L}
771 \DeclareInputText{"0E9}{\O}
772 \DeclareInputText{"0EA}{\OE}
773 \DeclareInputMath{"0EB}{\mathordmasculine}
774 \DeclareInputText{"0EC}{\@tabacckludge`o}
775 \DeclareInputText{"0ED}{\@tabacckludge'o}
776 \DeclareInputText{"0EE}{\^o}
777 \DeclareInputText{"0EF}{\~o}
778 \DeclareInputText{"0F0}{\"o}

779 \DeclareInputText{"0F1}{\ae}
780 \DeclareInputText{"0F2}{\@tabacckludge`u}
781 \DeclareInputText{"0F3}{\@tabacckludge'u}
782 \DeclareInputText{"0F4}{\^u}
783 \DeclareInputText{"0F5}{\i}
784 \DeclareInputText{"0F6}{\"u}
785 \DeclareInputText{"0F7}{\@tabacckludge'y}
786 \DeclareInputText{"0F8}{\l}
787 \DeclareInputText{"0F9}{\o}
788 \DeclareInputText{"0FA}{\oe}
789 \DeclareInputText{"0FB}{\ss}
790 %\DeclareInputText{"0FC}{?}
791 \DeclareInputText{"0FD}{\"y}

792 ⟨/next⟩

12 The Windows ANSI encoding

The coding for this input encoding was contributed by Berthold K.P. Horn (bkph@

 ai.mit.edu).

The Windows ANSI encoding file defines the characters in the Windows 3.1 ANSI encoding (Western Europe) — also known as code page 1252, which is based on ISO Latin 1 with important additions in the 128 - 159 range. Designed for: Danish, Dutch, English, Finnish, French, German, Icelandic, Italian, Norwegian, Portuguese, Spanish, and Swedish.

Note: Windows ANSI — like Macintosh standard roman encoding — has quotesingle at 39, and grave at 96 — which is here ignored.

793 ⟨∗ansinew⟩

794 \DeclareInputText{"082}{\quotesinglbase}
795 \DeclareInputText{"083}{\textflorin}
796 \DeclareInputText{"084}{\quotedblbase}
797 \DeclareInputText{"085}{\ldots}
798 \DeclareInputText{"086}{\dag}
799 \DeclareInputText{"087}{\ddag}
800 \DeclareInputText{"088}{\^{}}
801 \DeclareInputText{"089}{\perthousand}
802 \DeclareInputText{"08A}{\v S}
803 \DeclareInputText{"08B}{\guilsinglleft}
804 \DeclareInputText{"08C}{\OE}
805 \DeclareInputText{"091}{\textquoteleft}
806 \DeclareInputText{"092}{\textquoteright}
807 \DeclareInputText{"093}{\textquotedblleft}
808 \DeclareInputText{"094}{\textquotedblright}
809 \DeclareInputText{"095}{\textbullet}
810 \DeclareInputText{"096}{\textemdash}
811 \DeclareInputText{"097}{\textendash}
812 \DeclareInputText{"098}{\~{}}
813 \DeclareInputText{"099}{\texttrademark}
814 \DeclareInputText{"09A}{\v s}

815 \DeclareInputText{"09B}{\guilsinglright}
816 \DeclareInputText{"09C}{\oe}
817 \DeclareInputText{"09F}{\"y}
818 \DeclareInputText{"0A0}{\nobreakspace}
819 \DeclareInputMath{"0B0}{\mathdegree}
820 \DeclareInputText{"0A1}{\textexclamdown}
821 \DeclareInputMath{"0B1}{\pm}
822 \DeclareInputText{"0A2}{\textcent}
823 \DeclareInputMath{"0B2}{\mathtwosuperior}
824 \DeclareInputText{"0A3}{\pounds}
825 \DeclareInputMath{"0B3}{\maththreesuperior}
826 \DeclareInputText{"0A4}{\textcurrency}
827 \DeclareInputText{"0B4}{\@tabacckludge'{}}
828 \DeclareInputText{"0A5}{\textyen}
829 \DeclareInputMath{"0B5}{\mu}
830 \DeclareInputText{"0A6}{\textbrokenbar}
831 \DeclareInputText{"0B6}{\P}
832 \DeclareInputText{"0A7}{\S}
833 \DeclareInputText{"0B7}{\textperiodcentered}
834 \DeclareInputText{"0A8}{\"{}}
835 \DeclareInputText{"0B8}{\c\ }
836 \DeclareInputText{"0A9}{\copyright}
837 \DeclareInputMath{"0B9}{\mathonesuperior}
838 \DeclareInputMath{"0AA}{\mathordfeminine}
839 \DeclareInputMath{"0BA}{\mathordmasculine}
840 \DeclareInputText{"0AB}{\guillemotleft}
841 \DeclareInputText{"0BB}{\guillemotright}
842 \DeclareInputMath{"0AC}{\lnot}
843 \DeclareInputText{"0BC}{\textonequarter}
844 \DeclareInputText{"0AD}{\-}
845 \DeclareInputText{"0BD}{\textonehalf}
846 \DeclareInputText{"0AE}{\textregistered}
847 \DeclareInputText{"0BE}{\textthreequarters}
848 \DeclareInputText{"0AF}{\@tabacckludge={}}
849 \DeclareInputText{"0BF}{\textquestiondown}
850 \DeclareInputText{"0C0}{\@tabacckludge`A}
851 \DeclareInputText{"0D0}{\DH}
852 \DeclareInputText{"0C1}{\@tabacckludge'A}
853 \DeclareInputText{"0D1}{\~N}
854 \DeclareInputText{"0C2}{\^A}
855 \DeclareInputText{"0D2}{\@tabacckludge`O}
856 \DeclareInputText{"0C3}{\~A}
857 \DeclareInputText{"0D3}{\@tabacckludge'O}
858 \DeclareInputText{"0C4}{\"A}
859 \DeclareInputText{"0D4}{\^O}
860 \DeclareInputText{"0C5}{\AA}
861 \DeclareInputText{"0D5}{\~O}
862 \DeclareInputText{"0C6}{\AE}
863 \DeclareInputText{"0D6}{\"O}

864 \DeclareInputText{"0C7}{\c C}
865 \DeclareInputMath{"0D7}{\times}
866 \DeclareInputText{"0C8}{\@tabacckludge`E}
867 \DeclareInputText{"0D8}{\O}
868 \DeclareInputText{"0C9}{\@tabacckludge'E}
869 \DeclareInputText{"0D9}{\@tabacckludge`U}
870 \DeclareInputText{"0CA}{\^E}
871 \DeclareInputText{"0DA}{\@tabacckludge'U}
872 \DeclareInputText{"0CB}{\"E}
873 \DeclareInputText{"0DB}{\^U}
874 \DeclareInputText{"0CC}{\@tabacckludge`I}
875 \DeclareInputText{"0DC}{\"U}
876 \DeclareInputText{"0CD}{\@tabacckludge'I}
877 \DeclareInputText{"0DD}{\@tabacckludge'Y}
878 \DeclareInputText{"0CE}{\^I}
879 \DeclareInputText{"0DE}{\TH}
880 \DeclareInputText{"0CF}{\"I}
881 \DeclareInputText{"0DF}{\ss}
882 \DeclareInputText{"0E0}{\@tabacckludge`a}
883 \DeclareInputText{"0F0}{\dh}
884 \DeclareInputText{"0E1}{\@tabacckludge'a}
885 \DeclareInputText{"0F1}{\~n}
886 \DeclareInputText{"0E2}{\^a}
887 \DeclareInputText{"0F2}{\@tabacckludge`o}
888 \DeclareInputText{"0E3}{\~a}
889 \DeclareInputText{"0F3}{\@tabacckludge'o}
890 \DeclareInputText{"0E4}{\"a}
891 \DeclareInputText{"0F4}{\^o}
892 \DeclareInputText{"0E5}{\aa}
893 \DeclareInputText{"0F5}{\~o}
894 \DeclareInputText{"0E6}{\ae}
895 \DeclareInputText{"0F6}{\"o}
896 \DeclareInputText{"0E7}{\c c}
897 \DeclareInputMath{"0F7}{\div}
898 \DeclareInputText{"0E8}{\@tabacckludge`e}
899 \DeclareInputText{"0F8}{\o}
900 \DeclareInputText{"0E9}{\@tabacckludge'e}
901 \DeclareInputText{"0F9}{\@tabacckludge`u}
902 \DeclareInputText{"0EA}{\^e}
903 \DeclareInputText{"0FA}{\@tabacckludge'u}
904 \DeclareInputText{"0EB}{\"e}
905 \DeclareInputText{"0FB}{\^u}
906 \DeclareInputText{"0EC}{\@tabacckludge`\i}
907 \DeclareInputText{"0FC}{\"u}
908 \DeclareInputText{"0ED}{\@tabacckludge'\i}
909 \DeclareInputText{"0FD}{\@tabacckludge'y}
910 \DeclareInputText{"0EE}{\^\i}
911 \DeclareInputText{"0FE}{\th}
912 \DeclareInputText{"0EF}{\"\i}

913 \DeclareInputText{"0FF}{\"y}

914 ⟨/ansinew⟩